# Demystifying NCCL: An In-depth Analysis of GPU Communication Protocols and Algorithms

### v1

@@ IV-B2 IB Verbs Transport
As an optimization to improve bandwidth utilization and reduce congestion, the IB transport instantiates 2 logical channels per remote GPU and per NIC (controlled by the NCHANNELS_PER_NET_PEER parameter) by default. Each logical channel maintains its own ncclIbSendComm structure, embedding an independent bundle of InfiniBand QPs. During execution, the host-side network proxy alternates between the two sendComm handles when issuing ncclNet->isend() calls, thereby splitting traffic across the QP sets. This lightweight round-robin strategy increases the effective per-QP chunk size, introduces path diversity for ECMP-aware fabrics, and enhances overall interconnect efficiency—all without incurring additional GPU-side state or coordination overhead.
QP Layout
For every pair of ranks, the RDMA plugin establishes two reliable connection (RC) QPs, one in each direction. The forward QP is responsible for the bulk data stream: the proxy issues one or more RDMA_WRITE work requests that push user data directly into the peer buffer, and finally a zero-byte RDMA_WRITE_WITH_IMM. The immediate data field of this last request encodes the transfer size and is polled by the receiver to detect completion. The reverse QP carries only a tiny clear-to-send (CTS) message, which is a single RDMA_WRITE that advertises the remote buffer address, rkeys, and tag information. While the same functionality could theoretically be multiplexed on a single QP, separating the CTS onto its own channel isolates latency-critical control traffic from the bandwidth-hungry data stream, allowing the network to deliver it with minimal head-of-line blocking.
Local Flush with Loop-back RDMA_READ
V NCCL Collective Algorithms
V-A Overview of Algorithm and Protocol Support
@@ V-C Iterative Execution of NCCL Collectives
V-D Qualitative Algorithm Analysis

### v2

@@ IV-B2 IB Verbs Transport
As an optimization to improve bandwidth utilization and reduce congestion, the IB transport instantiates 2 logical channels per remote GPU and per NIC (controlled by the NCHANNELS_PER_NET_PEER parameter) by default. Each logical channel maintains its own ncclIbSendComm structure, embedding an independent bundle of InfiniBand QPs. During execution, the host-side network proxy alternates between the two sendComm handles when issuing ncclNet->isend() calls, thereby splitting traffic across the QP sets. This round-robin strategy increases the effective per-QP chunk size, introduces path diversity for ECMP-aware fabrics, and enhances overall interconnect efficiency—all without incurring additional coordination overhead.
@@
QP Layout
For every pair of ranks, the RDMA plugin establishes two reliable connection (RC) QPs, one in each direction. The forward QP is responsible for the bulk data stream: the proxy issues one or more RDMA_WRITE work requests that push user data directly into the peer buffer, followed by an RDMA_WRITE_WITH_IMM for completion notification. The behavior of this final request depends on the communication pattern: for small messages, it combines data and size information in a single operation to minimize overhead; for large messages with adaptive routing, it sends a zero-byte write with the transfer size as immediate data to guarantee ordering while allowing the bulk data to use adaptive routing; for aggregated multi-message transfers, it writes precise size information to a remote array while using immediate data for completion signaling. The reverse QP carries only a tiny clear-to-send (CTS) message, which is a single RDMA_WRITE that advertises the remote buffer address, rkeys, and tag information. While the same functionality could theoretically be multiplexed on a single QP, separating the CTS onto its own channel isolates latency-critical control traffic from the bandwidth-hungry data stream, allowing the network to deliver it with minimal head-of-line blocking.
Local Flush with Loop-back RDMA_READ
V NCCL Collective Algorithms
V-A Overview of Algorithm and Protocol Support
@@ V-C Iterative Execution of NCCL Collectives
V-D Mapping Communication Pipelines to CUDA Hierarchy
To fully understand NCCL’s performance characteristics, it is crucial to examine how these communication channels and data pipelines map onto the GPU’s parallel execution model. NCCL orchestrates hundreds of thousands of threads across multiple SMs through a carefully structured hierarchy that aligns with CUDA’s thread organization.
V-D1 Grid and Block Structure
NCCL kernels are launched with a grid dimension of (nChannels, 1, 1), where nChannels represents the number of active communication channels for the operation. This one-to-one mapping ensures that each CUDA block (blockIdx.x) corresponds to exactly one communication channel. Within each block, NCCL uses a variable number of threads, ranging from NCCL_MIN_NTHREADS to NCCL_MAX_NTHREADS. The exact thread count is determined by NCCL’s autotuning system and stored in plan->threadPerBlock during kernel launch preparation.
V-D2 Channel-to-Block ID Mapping
The mapping between blockIdx.x and channel ID is managed through a bitmask-based approach. Each kernel receives a channelMask that specifies which channels are active for the current operation. The device code computes the correspondence between block index and channel ID using bit-population counts: for a given blockIdx.x, the channel ID is determined by finding the blockIdx.x-th set bit in the channelMask.
V-D3 Warp-Level Organization
Within each block, NCCL organizes threads into specialized roles at the warp level. The first two warps handle initialization tasks: warp 0 loads the communicator metadata (ncclDevComm) into shared memory, while warp 1 loads the channel-specific data (ncclDevChannel). The remaining warps perform the actual communication and computation work.
For collective operations, the number of working warps is controlled by the nWarps field in the work descriptor. Different algorithms allocate these warps differently. For example, in NVLS AllReduce, warps are subdivided into groups for different phases. For point-to-point operations, warps are partitioned between send and receive operations, with the distribution computed dynamically based on the number of concurrent transfers.
V-D4 Slot-Based Pipeline Execution
Within each channel’s buffer, the NCCL_STEPS slots enable fine-grained pipelining at the thread level. Threads within a warp collaborate to move data through these slots in a circular manner. Each slot contains a ncclConnFifo structure with fields for mode, offset, size, and data pointer, allowing data to exist in different pipeline states simultaneously: being computed, queued for transmission, in-flight over the network, or ready for consumption.
V-D5 Thread-Level Data Movement
At the finest granularity, NCCL distributes work among individual threads within each warp. For bulk data movement, threads process multiple data elements per iteration in an unrolled fashion. The exact amount depends on the protocol.
Thread divergence is carefully managed through warp-uniform operations. For instance, all threads in a warp perform the same sequence of operations (send, reduce, copy) but on different data elements or memory addresses. This ensures maximum utilization of the GPU’s SIMT architecture.
V-D6 Concurrent Pipeline Execution
Critically, NCCL never executes ”one communication task” as might be suggested by high-level descriptions. Instead, it maintains dozens of concurrent data movement pipelines simultaneously. Multiple channels execute in parallel across different SMs, multiple slots within each channel pipeline different stages of data flow, and multiple warps within each channel handle different stages of the communication. This multi-level parallelism is essential for achieving high bandwidth utilization.
@@ V-D6 Concurrent Pipeline Execution
V-E Qualitative Algorithm Analysis